# Differentiable Pareto-Smoothed Weighting for High-Dimensional Heterogeneous Treatment Effect Estimation

### v1

@@ 3.2.2 Weight Replacement with GPD Quantiles
where $\hat{\operatorname{\mathrm{F}}}$ denotes the fitted GPD cumulative distribution function. By contrast, we do not change the other weight values, i.e., $w_{[1]},\dots,w_{[n-M]}$. Hence, letting $i=n-M+m$, we can summarize the weight replacement procedure as
@@
In this paper, we propose to utilize the weight replacement formula in (21) to improve the estimation stability of weighted representation learning. Unfortunately, we cannot directly employ this formula in an end-to-end manner because it needs non-differentiable computations.
@@ 3.3 Non-Differentiable Procedures
Ranking is an operation that takes input vector $\textbf{{w}}=[w_{1},\dots,w_{n}]^{\top}$ and outputs the position of each element $w_{i}$ in the sorted vector $[w_{[1]},\dots,w_{[n]}]^{\top}$, where $w_{[1]}\leq\dots\leq w_{[n]}$. To illustrate this operation, consider the case with $n=3$. For instance, if w satisfies $w_{3}\leq w_{1}\leq w_{2}$, since $w_{1}=w_{[2]}$, $w_{2}=w_{[3]}$, and $w_{3}=w_{[1]}$ hold, the rank of w is given as vector $\textbf{{r}}=[2,3,1]^{\top}$. Formally, such an operation can be expressed as $\textbf{{r}}=r(\textbf{{w}})$, using function $r$, called a rank function (See Appendix A for the definition of function $r$).
Unfortunately, this rank function is not differentiable with respect to input w. To see this, consider $\textbf{{w}}=[w_{1},1,2,3]^{\top}$ and observe how the rank of $w_{1}$ varies when we increase $w_{1}$’s value. In this case, its rank, $r_{1}$, is given as a piecewise constant function, as illustrated as the black line in Figure 2. Since the derivative of such a piecewise constant function is always zero or undefined, we cannot perform gradient backpropagation and hence cannot employ the weight correction technique in (21) in an end-to-end manner. Therefore, with such a non-differentiable rank function, we cannot use Pareto smoothing for weighted representation learning, which jointly learns the propensity score model, the feature representations, and the outcome prediction models.
Figure 2: Illustration of rank function $\textbf{{r}}=r(\textbf{{w}})$ (black) and differentiable rank functions $\textbf{{r}}=r_{\varepsilon}(\textbf{{w}})$ (orange and green). Here we take input vector $\textbf{{w}}=[w_{1},1,2,3]^{\top}$, vary $w_{1}$’s value, and look at how its rank $_{1}\in\textbf{{r}}$ r changes. When regularization parameter $\varepsilon\rightarrow 0$, $r_{\varepsilon}$ converges to $r$ [Blondel et al., 2020].
One may consider a separate learning approach that trains the propensity score model, computes the Pareto-smoothed IPW weights by (21), and then learns the feature representations. This approach, however, requires to fit a propensity score model directly to features X, not their representation. Since accurately estimating a propensity score model from high-dimensional features X is considerably difficult, due to the model misspecification error, such a separate learning approach will lead to CATE estimation bias. We experimentally show its poor performance in Section 4.1.
@@ 3.4.1 Differentiable Approximation
The weight replacement formula in (21) requires the computation of the two troublesome piecewise constant functions. One is rank function $r$, which is needed to obtain the position of weight $w_{i}$ in the sorted vector $[w_{[1]},\dots,w_{[n]}]^{\top}$, and the other is indicator function $\operatorname{\mathbf{I}}(i\geq n-M+1)$.
@@
Regarding indicator function $\operatorname{\mathbf{I}}$ in (21), we approximate it as the sigmoid function $\varsigma$:
@@ 3.4.2 Reformulation of GPD Parameter Estimators
As regards location parameter $\hat{\mu}$ in (14), since this estimator is given as $w_{[n-M]}$, i.e., the largest weight among $w_{[1]},\dots,w_{[n-M]}$, we reformulate it as
@@ 3.4.3 Overall Algorithm
Compared with DRCFR [Hassanpour and Greiner, 2020], our method requires additional time to compute Pareto-smoothed weights (lines 12-16 in Algorithm 1). In particular, computing differentiable rank (line 12) requires the time complexity $\mathcal{O}(B\mathrm{log}B)$ for mini-batch size $B$, which is needed to evaluate the objective function in (3) and its gradient for each iteration in the training phase.
Remark: Strictly speaking, the choice of activation functions in propensity score $\pi$ and feature representations $\Gamma$ and $\Delta$ is important to satisfy the assumption of Pareto smoothing that the distribution of the importance sampling weight is absolutely continuous, which is employed to prove the asymptotic consistency in Theorem 1 of Vehtari et al. [2024]. This assumption holds if each activation is differentiable almost everywhere (i.e., differentiable except on a set of measure zero). However, for instance, using the rectified linear unit (ReLU) in propensity score model $\pi$ makes the distribution of IPW weight discontinuous, thus violating the assumption of Pareto smoothing.
@@ 4.1 Semi-synthetic Data
The News dataset is constructed from $n=5000$ articles, randomly sampled from the New York Times corpus in the UCI repository.  [^3] https://archive.ics.uci.edu/dataset/164/bag+of+words The task is to infer the effect of the viewing device (desktop ($A=0$) or mobile ($A=1$)) on the readers’ experience $Y$. Features X are the count of $d=3477$ words in each article. Treatment $A$ and outcome $Y$ are simulated using the latent topic variables obtained by fitting a topic model on X. The ACIC dataset is derived from the clinical measurements of $d=177$ features in the Linked Birth and Infant Death Data (LBIDD) [MacDorman and Atkinson, 1998] and is developed for the data analysis competition called ACIC2018. Here, we randomly picked up $n=5000$ observations and prepared $20$ datasets. With both semi-synthetic datasets, we randomly split each sample into training, validation, and test data with a ratio of 60/20/20.
@@
Baselines: To evaluate our method (DPSW) and its variant that performs self-normalization (DPSW Norm.), we consider $15$ baselines. With DRCFR [Hassanpour and Greiner, 2020], we tested the four different weighting schemes: no weight modification (DRCFR), self-normalization (DRCFR Norm.; Eq. (9)), the weight truncation with the threshold suggested by Crump et al. [2009] (DRCFR Trunc.; Eq. (8)), and the scheme that ignores the prediction loss for the individuals with extreme weights based on the same threshold (DRCFR Ignore). We also tested a separate learning approach (PSW; Section 3.3), which trains propensity score $\pi$ with $\{(a_{i},\textbf{{x}}_{i})\}_{i=1}^{n}$ beforehand and learns only $\Delta(\textbf{{X}})$ and $\Upsilon(\textbf{{X}})$ using the Pareto-smoothed IPW weights. Other baselines include (i) linear regression methods: a single model with treatment $A$ added to its input (LR-1) and two separate models for each treatment (LR-2); (ii) meta-learner methods: the S-Learner (SL), the T-Learner (TL), the X-Learner (XL), and the DR-Learner (DRL); (iii) tree-based methods: the causal forest [Athey et al., 2019] (CF) and the variant combined with double/debiased machine learning [Chernozhukov et al., 2018] (CF DML); and (iv) neural network methods: the treatment-agnostic regression network [Shalit et al., 2017] (TARNet) and the generative adversarial network [Yoon et al., 2018] (GANITE). We detail the settings of these baselines in Section B.1.
@@
Settings: Regarding our method and DRCFR, we used three-layered feed-forward neural networks (FNNs) to formulate feature representations $\Gamma(\textbf{{X}})$, $\Delta(\textbf{{X}})$, and $\Upsilon(\textbf{{X}})$, propensity score $\pi$ and outcome prediction models $h^{0}$ and $h^{1}$.
@@
We tuned the hyperparameters (e.g., parameter $\varepsilon$ of differentiable rank $r_{\varepsilon}(\textbf{{w}})$ in our method) by minimizing the objective function value on validation data; such hyperparameter tuning is standard for CATE estimation [Shalit et al., 2017].
@@
Performance metric: Following Hill [2011], we used the precision in the estimation of heterogeneous effect (PEHE), $\mathrm{PEHE}\coloneqq\sqrt{\frac{1}{n}\left((y^{1}_{i}-y^{0}_{i})-\hat{\tau}_{i}\right)^{2}}$, where $y^{0}_{i}$ and $y^{1}_{i}$ are the true potential outcomes, and $\hat{\tau}_{i}$ denotes the predicted CATE value. We computed the mean and the standard deviation of test PEHE over $50$ realizations of potential outcomes (the News dataset) and $20$ realizations (the ACIC dataset).
@@
Our proposed framework (DPSW and DPSW Norm.) outperformed all baselines, demonstrating their effectiveness in CATE estimation from high-dimensional data. DPSW Norm. achieved lower PEHEs than DPSW, implying that the self-normalization of Pareto-smoothed weights can further improve the stability of the weight estimation.
Weighted representation learning methods (DR-CFR and DPSW) performed better than other neural network methods (TARNet and GANITE), especially on the ACIC dataset. Given that treatment $A$ and outcome $Y$ of this dataset were simulated using different features in X, these results emphasize the importance of performing data-driven feature separation via weighted representation in such a setup.
PSW worked much worse on the News dataset than DPSW, indicating that fitting a propensity score directly to high-dimensional features X leads to a severe model misspecification error, making the subsequent weighted representation learning difficult, even with the Pareto-smoothed weights. By contrast, our joint learning approach well performed, thanks to the use of differentiable Pareto smoothing.
@@ 4.2 Synthetic Data
Next, we investigated how well our method can learn the feature representations using synthetic data, where the data-generating processes are entirely known.
Figure 3: Learned encoder parameter differences and test PEHEs on synthetic data. (a): Value difference of $\mathbf{W}^{1}$ in encoder $\Gamma(\textbf{{X}})$; (b): Value difference of $\mathbf{W}^{1}$ in encoder $\Delta(\textbf{{X}})$; (c): Value difference of $\mathbf{W}^{1}$ in encoder $\Upsilon(\textbf{{X}})$; (d) Test PEHEs. With TARNet, since it learns single encoder, we computed all parameter value differences with weight matrix in same encoder.
Data: Following Hassanpour and Greiner [2020], we simulated synthetic data. We randomly generated features $\textbf{{X}}=[\textbf{{X}}_{\Gamma},\textbf{{X}}_{\Delta},\textbf{{X}}_{\Upsilon}]^{\top}\in\operatorname{\mathbb{R}}^{d}$ ($d=15,18,\dots,30$). Then, by regarding feature subsets $\textbf{{X}}_{\Gamma}\in\operatorname{\mathbb{R}}^{d/3}$, $\textbf{{X}}_{\Delta}\in\operatorname{\mathbb{R}}^{d/3}$, and $\textbf{{X}}_{\Upsilon}\in\operatorname{\mathbb{R}}^{d/3}$ as instrumental variables, confounders, and adjustment variables, respectively, we sampled binary treatment $A$ using $\textbf{{X}}_{\Gamma}$ and $\textbf{{X}}_{\Delta}$ and outcome $Y$ by employing $\textbf{{X}}_{\Delta}$ and $\textbf{{X}}_{\Upsilon}$ (See Section B.2 for the detail). We split each of $20$ datasets ($n=20000$) with a 50/25/25 training/validation/test ratio.
@@
where $\nu$ is exponential linear units (ELUs) [Clevert et al., 2016], and $\textbf{W}^{1}$, $\textbf{W}^{2}$, and $\textbf{W}^{3}$ are the weight parameter matrices in the first, the second, and the third layer, respectively.
To determine whether the learned FNN encoders correctly look at important features, we measured the attribution of features $\textbf{{X}}_{\Gamma}$, $\textbf{{X}}_{\Delta}$, and $\textbf{{X}}_{\Upsilon}$ by employing $\textbf{W}^{1}$, i.e., the trained weight matrix in the first layer of each encoder. For instance, we quantified $\textbf{{X}}_{\Gamma}$’s attribution on $\Gamma(\textbf{{X}})$, by taking two steps. First, we partitioned its learned weight parameter matrix as $\textbf{W}^{1}=[\textbf{W}^{1}_{\Gamma},\textbf{W}^{1}_{-\Gamma}]$, where $\textbf{W}^{1}_{\Gamma}$ is the submatrix with the first $d/3$ columns of $\textbf{W}^{1}$ and $\textbf{W}^{1}_{-\Gamma}$ is the one with the other columns. Then we measured how greatly features $\textbf{{X}}_{\Gamma}$ affects the learned representation $\Gamma(\textbf{{X}})$ by taking the relative difference between the average absolute values of the weight parameter submatrices, i.e., $\frac{\overline{|\textbf{W}^{1}_{\Gamma}|}-\overline{|\textbf{W}^{1}_{-\Gamma}|}}{\overline{|\textbf{W}^{1}_{-\Gamma}|}}$. We evaluated other learned representations, $\Delta(\textbf{{X}})$ and $\Upsilon(\textbf{{X}})$, in the same way.
Results: Figure 3 shows the means and standard deviations of learned parameter differences and test PEHEs over $20$ randomly generated synthetic datasets.
With our DPSW method and DRCFR, the absolute parameter values $|\textbf{W}^{1}_{\Gamma}|$, $|\textbf{W}^{1}_{\Delta}|$, and $|\textbf{W}^{1}_{\Gamma}|$ were sufficiently larger than $|\textbf{W}^{1}_{-\Gamma}|$, $|\textbf{W}^{1}_{-\Delta}|$, and $|\textbf{W}^{1}_{-\Gamma}|$, respectively, showing that both methods correctly learn $\Gamma(\textbf{{X}})$, $\Delta(\textbf{{X}})$, and $\Upsilon(\textbf{{X}})$ that are highly dependent on instrumental variables $\textbf{{X}}_{\Gamma}$, confounders $\textbf{{X}}_{\Delta}$, and adjustment variables $\textbf{{X}}_{\Upsilon}$, respectively. These results made a clear contrast to TARNet, which learns a single representation without performing feature separation.
The same is true for the CATE estimation performance (Figure 3 (d)). The test PEHE of TARNet was larger than DPSW and DRCFR, demonstrating the importance of data-driven feature separation via weighted representation learning. By contrast, our method achieved the lowest PEHE, thus indicating that our weight correction framework can successfully improve the CATE estimation performance of DRCFR.
Performance under high-dimensional setup: We confirmed that our method also worked well with $d=600,1200,\dots,3000$ (See Appendix C for the detail).

### v2

@@ 3.2.2 Weight Replacement with GPD Quantiles
where $\hat{\operatorname{\mathrm{F}}}$ denotes a fitted GPD cumulative distribution function. By contrast, we do not change the other weight values, i.e., $w_{[1]},\dots,w_{[n-M]}$. Hence, letting $i=n-M+m$, we can summarize the weight replacement procedure as
@@
In this paper, we utilize the weight replacement formula in (21) to improve the estimation stability of weighted representation learning. Unfortunately, we cannot directly employ this formula in an end-to-end manner because it needs non-differentiable computations.
@@ 3.3 Non-Differentiable Procedures
Ranking is an operation that takes input vector $\textbf{{w}}=[w_{1},\dots,w_{n}]^{\top}$ and outputs the position of each element $w_{i}$ in sorted vector $[w_{[1]},\dots,w_{[n]}]^{\top}$, where $w_{[1]}\leq\dots\leq w_{[n]}$. To illustrate this operation, consider a case with $n=3$. For instance, if w satisfies $w_{3}\leq w_{1}\leq w_{2}$, since $w_{1}=w_{[2]}$, $w_{2}=w_{[3]}$, and $w_{3}=w_{[1]}$ hold, the rank of w is given as vector $\textbf{{r}}=[2,3,1]^{\top}$. Formally, such an operation can be expressed as $\textbf{{r}}=r(\textbf{{w}})$, using function $r$, called a rank function (See Appendix A for the definition of function $r$).
Unfortunately, this rank function is not differentiable with respect to input w. To see this, consider $\textbf{{w}}=[w_{1},1,2,3]^{\top}$ and observe how the rank of $w_{1}$ varies when we increase its value. In this case, its rank, $r_{1}$, is given as a piecewise constant function, as illustrated as the black line in Figure 2. Since the derivative of such a piecewise constant function is always zero or undefined, we cannot perform gradient backpropagation and hence cannot employ the weight correction technique in (21) in an end-to-end manner. Therefore, with such a non-differentiable rank function, we cannot use Pareto smoothing for weighted representation learning, which jointly learns the propensity score model, the feature representations, and the outcome prediction models.
Figure 2: Illustration of rank function $\textbf{{r}}=r(\textbf{{w}})$ (black) and differentiable rank functions $\textbf{{r}}=r_{\varepsilon}(\textbf{{w}})$ (orange and green): Here we take input vector $\textbf{{w}}=[w_{1},1,2,3]^{\top}$, vary $w_{1}$’s value and look at how its rank $_{1}\in\textbf{{r}}$ r changes. When regularization parameter $\varepsilon\rightarrow 0$, $r_{\varepsilon}$ converges to $r$ [Blondel et al., 2020].
One may consider a separate learning approach that trains the propensity score model, computes the Pareto-smoothed IPW weights by (21), and learns the feature representations. This approach, however, requires directly fitting a propensity score model to features X, not their representations. Since accurately estimating a propensity score model from high-dimensional features X is considerably difficult, due to the model misspecification error, such a separate learning approach leads to CATE estimation bias. We experimentally show its poor performance in Section 4.1.
@@ 3.4.1 Differentiable Approximation
The weight replacement formula in (21) requires the computation of two troublesome piecewise constant functions. One is rank function $r$, which is needed to obtain the position of weight $w_{i}$ in sorted vector $[w_{[1]},\dots,w_{[n]}]^{\top}$, and the other is indicator function $\operatorname{\mathbf{I}}(i\geq n-M+1)$.
@@
As a differentiable approximation of indicator function $\operatorname{\mathbf{I}}$ in (21), we employ sigmoid function $\varsigma$:
@@ 3.4.2 Reformulation of GPD Parameter Estimators
Regarding location parameter $\hat{\mu}$ in (14), since this estimator is given as $w_{[n-M]}$, i.e., the largest weight among $w_{[1]},\dots,w_{[n-M]}$, we reformulate it as
@@ 3.4.3 Overall Algorithm
Compared with DRCFR [Hassanpour and Greiner, 2020], our method requires additional time to compute Pareto-smoothed weights (lines 12-16 in Algorithm 1). In particular, computing differentiable rank (line 12) requires time complexity $\mathcal{O}(B\mathrm{log}B)$ for mini-batch size $B$, which is needed to evaluate the objective function in (3) and its gradient for each iteration in the training phase.
Remark: Strictly speaking, the choice of activation functions in propensity score $\pi$ and feature representations $\Gamma$ and $\Delta$ is critical for satisfying the assumption of Pareto smoothing that the distribution of the importance sampling weight is absolutely continuous, which is employed to prove the asymptotic consistency in Theorem 1 of Vehtari et al. [2024]. This assumption holds if each activation is differentiable almost everywhere (i.e., differentiable except on a set of measure zero). However, for instance, using the rectified linear unit (ReLU) in propensity score model $\pi$ makes the distribution of IPW weight discontinuous, thus violating the assumption of Pareto smoothing. Even with almost everywhere differentiable activation functions, due to the lack of learning theory on neural network models, deriving the asymptotic consistency of our CATE estimator is extremely challenging and is left as our future work.
@@ 4.1 Semi-synthetic Data
The News dataset is constructed from $n=5000$ articles, randomly sampled from the New York Times corpus in the UCI repository.  [^3] https://archive.ics.uci.edu/dataset/164/bag+of+words The task is to infer the effect of the viewing device (desktop ($A=0$) or mobile ($A=1$)) on the readers’ experience $Y$. Features X are the count of $d=3477$ words in each article. Treatment $A$ and outcome $Y$ are simulated using the latent topic variables obtained by fitting a topic model on X. The ACIC dataset is derived from the clinical measurements of $d=177$ features in the Linked Birth and Infant Death Data (LBIDD) [MacDorman and Atkinson, 1998], which was developed for a data analysis competition called ACIC2018. We randomly selected $n=5000$ observations and prepared $20$ datasets. With both semi-synthetic datasets, we randomly split each sample into training, validation, and test data with a ratio of 60/20/20.
@@
Baselines: To evaluate our method (DPSW) and its variant that performs self-normalization (DPSW Norm.), we consider $15$ baselines. With DRCFR [Hassanpour and Greiner, 2020], we tested four different weighting schemes: no weight modification (DRCFR), self-normalization (DRCFR Norm.; Eq. (9)), weight truncation with a threshold suggested by Crump et al. [2009] (DRCFR Trunc.; Eq. (8)), and a scheme that ignores the prediction loss for individuals with extreme weights based on the same threshold (DRCFR Ignore). We also tested a separate learning approach (PSW; Section 3.3), which trains propensity score $\pi$ with $\{(a_{i},\textbf{{x}}_{i})\}_{i=1}^{n}$ beforehand and learns only $\Delta(\textbf{{X}})$ and $\Upsilon(\textbf{{X}})$ using Pareto-smoothed IPW weights. Other baselines include (i) linear regression methods: a single model with treatment $A$ added to its input (LR-1) and two separate models for each treatment (LR-2); (ii) meta-learner methods: the S-Learner (SL), the T-Learner (TL), the X-Learner (XL), and the DR-Learner (DRL); (iii) tree-based methods: causal forest [Athey et al., 2019] (CF) and a variant combined with double/debiased machine learning [Chernozhukov et al., 2018] (CF DML); and (iv) neural network methods: the treatment-agnostic regression network [Shalit et al., 2017] (TARNet) and the generative adversarial network [Yoon et al., 2018] (GANITE). We describe the settings of these baselines in Section B.1.
@@
Settings: Regarding our method and DRCFR, we used three-layered feed-forward neural networks (FNNs) to formulate feature representations $\Gamma(\textbf{{X}})$, $\Delta(\textbf{{X}})$, and $\Upsilon(\textbf{{X}})$, propensity score $\pi$, and outcome prediction models $h^{0}$ and $h^{1}$.
@@
We tuned the hyperparameters (e.g., parameter $\varepsilon$ of differentiable rank $r_{\varepsilon}(\textbf{{w}})$ in our method) by minimizing the objective function value on the validation data; such hyperparameter tuning is standard for CATE estimation [Shalit et al., 2017].
@@
Performance metric: Following Hill [2011], we used the precision in the estimation of heterogeneous effect (PEHE), $\mathrm{PEHE}\coloneqq\sqrt{\frac{1}{n}\left((y^{1}_{i}-y^{0}_{i})-\hat{\tau}_{i}\right)^{2}}$, where $y^{0}_{i}$ and $y^{1}_{i}$ are the true potential outcomes, and $\hat{\tau}_{i}$ denotes the predicted CATE value. We computed the mean and the standard deviation of the test PEHEs over $50$ realizations of potential outcomes (the News dataset) and $20$ realizations (the ACIC dataset).
@@
Our proposed frameworks (DPSW and DPSW Norm.) outperformed all the baselines, demonstrating their effectiveness in CATE estimation from high-dimensional data. DPSW Norm. achieved lower PEHEs than DPSW, implying that the self-normalization of Pareto-smoothed weights further improves the stability of the weight estimation.
Weighted representation learning methods (DR-CFR and DPSW) outperformed the other neural network methods (TARNet and GANITE), especially on the ACIC dataset. Given that treatment $A$ and outcome $Y$ of this dataset were simulated using different features in X, these results emphasize the importance of performing data-driven feature separation by weighted representation in such a setup.
PSW worked much worse on the News dataset than DPSW, indicating that fitting a propensity score directly to high-dimensional features X leads to severe model misspecification error, complicating subsequent weighted representation learning, even with Pareto-smoothed weights. By contrast, our joint learning approach performed well, since it used differentiable Pareto smoothing.
@@ 4.2 Synthetic Data
Next we investigated how well our method learned the feature representations using synthetic data, where the data-generating processes are entirely known.
Figure 3: Learned encoder parameter differences and test PEHEs on synthetic data: (a): value difference of $\mathbf{W}^{1}$ in encoder $\Gamma(\textbf{{X}})$; (b): value difference of $\mathbf{W}^{1}$ in encoder $\Delta(\textbf{{X}})$; (c): value difference of $\mathbf{W}^{1}$ in encoder $\Upsilon(\textbf{{X}})$; (d) test PEHEs. With TARNet, since it learns a single encoder, we computed all parameter value differences with weight matrix in same encoder.
Data: Following Hassanpour and Greiner [2020], we simulated the synthetic data. We randomly generated features $\textbf{{X}}=[\textbf{{X}}_{\Gamma},\textbf{{X}}_{\Delta},\textbf{{X}}_{\Upsilon}]^{\top}\in\operatorname{\mathbb{R}}^{d}$ ($d=15,18,\dots,30$). Next, by regarding feature subsets $\textbf{{X}}_{\Gamma}\in\operatorname{\mathbb{R}}^{d/3}$, $\textbf{{X}}_{\Delta}\in\operatorname{\mathbb{R}}^{d/3}$, and $\textbf{{X}}_{\Upsilon}\in\operatorname{\mathbb{R}}^{d/3}$ as instrumental variables, confounders, and adjustment variables, respectively, we sampled binary treatment $A$ using $\textbf{{X}}_{\Gamma}$ and $\textbf{{X}}_{\Delta}$ and outcome $Y$ by employing $\textbf{{X}}_{\Delta}$ and $\textbf{{X}}_{\Upsilon}$ (See Section B.2 for details). We split each of the $20$ datasets ($n=20000$) with a 50/25/25 training/validation/test ratio.
@@
where $\nu$ is the exponential linear units (ELUs) [Clevert et al., 2016], and $\textbf{W}^{1}$, $\textbf{W}^{2}$, and $\textbf{W}^{3}$ are the weight parameter matrices in the first, second, and third layer, respectively.
To determine whether the learned FNN encoders correctly look at important features, we measured the attribution of features $\textbf{{X}}_{\Gamma}$, $\textbf{{X}}_{\Delta}$, and $\textbf{{X}}_{\Upsilon}$ by employing $\textbf{W}^{1}$, i.e., the trained weight matrix in the first layer of each encoder. For instance, we quantified $\textbf{{X}}_{\Gamma}$’s attribution on $\Gamma(\textbf{{X}})$ in two steps. First, we partitioned its learned weight parameter matrix as $\textbf{W}^{1}=[\textbf{W}^{1}_{\Gamma},\textbf{W}^{1}_{-\Gamma}]$, where $\textbf{W}^{1}_{\Gamma}$ is a submatrix with the first $d/3$ columns of $\textbf{W}^{1}$ and $\textbf{W}^{1}_{-\Gamma}$ is the one with the other columns. Then we measured how greatly features $\textbf{{X}}_{\Gamma}$ affect learned representation $\Gamma(\textbf{{X}})$ by taking the relative difference between the average absolute values of the weight parameter submatrices, i.e., $\frac{\overline{|\textbf{W}^{1}_{\Gamma}|}-\overline{|\textbf{W}^{1}_{-\Gamma}|}}{\overline{|\textbf{W}^{1}_{-\Gamma}|}}$. We evaluated other learned representations, $\Delta(\textbf{{X}})$ and $\Upsilon(\textbf{{X}})$, in the same way.
Results: Figure 3 shows the means and standard deviations of the learned parameter differences and the test PEHEs over $20$ randomly generated synthetic datasets.
With our DPSW method and DRCFR, absolute parameter values $|\textbf{W}^{1}_{\Gamma}|$, $|\textbf{W}^{1}_{\Delta}|$, and $|\textbf{W}^{1}_{\Gamma}|$ were sufficiently larger than $|\textbf{W}^{1}_{-\Gamma}|$, $|\textbf{W}^{1}_{-\Delta}|$, and $|\textbf{W}^{1}_{-\Gamma}|$, respectively, showing that both methods correctly learned $\Gamma(\textbf{{X}})$, $\Delta(\textbf{{X}})$, and $\Upsilon(\textbf{{X}})$ that are highly dependent on instrumental variables $\textbf{{X}}_{\Gamma}$, confounders $\textbf{{X}}_{\Delta}$, and adjustment variables $\textbf{{X}}_{\Upsilon}$, respectively. These results offer a clear contrast to TARNet, which learns a single representation without feature separation.
The same is true for the CATE estimation performance (Figure 3 (d)). TARNet’s test PEHE was larger than DPSW and DRCFR, demonstrating the importance of data-driven feature separation by weighted representation learning. By contrast, our method achieved the lowest PEHE, thus indicating that our weight correction framework successfully improved the CATE estimation performance of DRCFR.
Performance under high-dimensional setup: We confirmed that our method also worked well with $d=600,1200,\dots,3000$ (See Appendix C for details).